Exercise 7. Let X be a standard normal random variable. Prove that for any integer n > 0, ELY?] = 1207) and E[x2n+] = 0.
Answer #1

# Let x be standard normal distrebution, drea I pdf of x is given by f(x) = 1 / 2 XER. ( E(X) = so can fexs de I de 8 - 1 x²Elxan) = 2h 5 that et at = 2 Penth) nt where p is gamma function We know that Tez T (2tb) = 21-22 J r (22) Therefore Pentt) =(i) E (xant y = fm sent for de ant ant ant E Cx2ht & o Integrand is odd function.